OpenStudy (anonymous):

the ratio of the radii of two spheres is 6 : 5, what is the ratio of the surface areas of the two spheres?

Directrix (directrix):

Theorem: If two solids are similar, the square of the scale factor of the two solids is equal to the ratio of any two corresponding area measurements of the solids.

Directrix (directrix):

All spheres are similar. So, according to the theorem, the ratio of the surface areas of the two spheres is the following: (6/5)^2 = (6/5) * (6/5) = ? @spaghettigoo123